Two numbers are in the ratio 5:3. If they differ by 18, what are the numbers? 

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Let the two numbers be 5x and 3x. According to the question, 5x-3x=18
⇒2x=18
⇒x=18/2
⇒x=19
Thus,
The numbers are 5x = 5 × 9 = 45
And3x=3×9=27.

Thanks